### v1

@@ Abstract
Keywords:
Attack graph Attack modelling technique Directed acyclic graph Attack surface Multilayer networkLateral movement
@@ 1 Introduction
Understanding the nature of cyber threats and protecting systems from attacks has become increasingly critical. Cyberattacks occur when malicious actors compromise the confidentiality, integrity, or availability of computer networks and/or services, by exploiting vulnerabilities therein. li2021comprehensive; wachter2023graph Cyber-attackers typically gain access to computer networks by using a series of exploits, in which each exploit facilitates the next. Once they have an initial entry point, they proceed with lateral movement to spread to the rest of the network and reach other hosts and applications within an organization. SMILIOTOPOULOS2024e26317 This series of exploits constitutes an attack path, while the collection of potential attack paths is an attack graph. zenitani2023attack; lallie2020review; zeng2019survey
@@
To improve the planning for network security, one needs first to assess the weaknesses in the system. The research presented in lof2010approach; CySeMoL introduced a model-based framework and a tool to analyse network security across different architectural scenarios. This framework uses a probabilistic model to represent attack paths and countermeasures, and it enables calculation of the probability of success of the attack based on the specific architectural scenario. A study in lippmann2006validating investigated strategies for protecting critical resources within enterprise networks, focusing mainly on the use of multiple layers of firewalls. This method deploys a perimeter firewall that separates the corporate network from the Internet, along with internal firewalls for critical subnets.
@@
In the present study, we investigate cyber-attacks using a novel approach that combines network modelling and graphs. noel2018review; lagraa2024review; idika2010extending This results in an attack graph that is a visual representation of potential paths an attacker could take to compromise a system. ou2011quantitative; stergiopoulos2022automatic; gain2023attack It describes the structure of attack paths as a directed acyclic graph (DAG) of nodes and links between them. kordy2014dag Each node in the attack graph can represent a host, network device, or vulnerability of a network system. zeng2019survey; Ou2011
@@
There are two commonly used types of attack graphs. The first one is a directed graph in which nodes denote system states related to security, such as file modifications, levels of user privileges, and trust relationships among hosts. The edges represent possible exploits or actions by the perpetrators. The second type involves nodes that denote the preconditions or postconditions of an exploit, with edges showing the effects of those conditions. Although our primary focus is on directed acyclic attack graphs, our model can also be used to include more general cyclic and recurrent graphs, or causal graphs. kuikkaSciRep
@@
An attack surface of a system refers to the various ways an intruder can exploit its weaknesses to gain access, ou2011quantitative including resources that a perpetrator can access. The more resources an attacker can access, the larger the attack surface, making the system less secure. To evaluate the security of a system based on its attack surface, several metrics have been developed. AttackSurfaceMetric
@@
In the present study, we introduce a quantitative attack modelling approach to combine vulnerability metrics of an enterprise network by integrating the attack graph and the structure of the communication network. Our probabilistic approach uses exploitability metrics to evaluate the likelihood that individual vulnerabilities will be exploited along the attack paths. lof2010approach; wang2008attack The model estimates the overall probability that a perpetrator successfully gains specific privileges or executes attacks. These probabilities are determined by merging alternative joining attack paths using probability theory for non-exclusive events. kuikkaSciRep
@@
This approach is novel, as it employs a detailed-level network modelling technique kuikkaSciRep that integrates the communication network and attack graph structures. It enables the construction of a quantitative model that describes cyber threats and attacks in the communication network. The outcomes of the model can be presented at varying levels of detail, incorporating attack graph nodes that represent microservices, communication network nodes like servers and network equipment, as well as user services and applications.
@@
Our approach is based on a framework of multilayer network models bianconi2018multilayer; interdonato2020multilayer and allows us to analyse cyber threats and design protective solutions against cyber attacks in enterprise networks. Here, we focus on investigating the infrastructure and communication layer, the cyber attack layer, and the application layer, each of which can be divided into sublayers. With the present approach, we can also evaluate the operation and robustness of the system from the point of view of physical network, protocol layers, cyber incidents, user services, and applications. All of these layers have the potential to simultaneously exert detrimental impacts on system users. Moreover, we can construct an extra layer for each user service to evaluate metrics related to both incoming and outgoing effects. This in turn allows us to analyse cyber threats targeting specific services or applications as well as the way an attack propagates within a service between microservices. Different services or application servers provide different options for detection, which affects the likelihood of successful mitigation. It is an essential feature of our method that individual attack paths and attack graphs can be extended to different services and servers in the network topology.
@@ 2 Related Work
In this section, we present a brief overview of related research on modelling cyberattacks and analysing their impact, risk, and resilience in enterprise networks. Here, we focus specifically on graph-based models, attack propagation models, and multilayer network models as well as the metrics to analyse network security. wang2017network; homer2013aggregating; pendleton2016survey In homer2013aggregating the authors present a probabilistic metric for network security by aggregating individual vulnerability measures derived from the scores of the Common Vulnerability Scoring System Scoring; wang2008attack (CVSS) in total network score. The survey on system security metrics presented in pendleton2016survey proposes a framework to measure system-level security by using the following four sub-metrics: system vulnerabilities, defense capabilities, attack severity, and situational factors. Here the authors introduce a hierarchical ontology to evaluate relationships among these sub-metrics. They also review existing security metrics and highlight their advantages and limitations to guide the development of more effective security metrics. The study in Noel2017 introduces a set of normalised metrics for comparisons across enterprises. It categorises related metrics into four families (victimization, network size, containment, and topology) that aggregate individual scores into family scores. These metrics are presented through interactive visualisations that show trends over time at individual, family, and overall network levels.
@@
In order to analyse a particular part of the network resilience, we need to apply network modelling techniques and graph models to characterise cyber attacks. noel2018review; lagraa2024review The probability formulas that are used to describe connectivity are also useful in explaining the concept of resilience. Ball_etal; kuikka2022resilience Here, the connectivity between two nodes is defined as the probability that there is a functioning connection between them. If the connectivity values between all pairs of neighbouring nodes in the network are known, the connectivity probabilities between any pair of nodes in the network can be calculated. Ball_etal This calculation can be performed using either simulation methods or algorithmic techniques. An alternative method to this was introduced in Almiala_Aalto_Kuikka, where probabilities were calculated based on spreading probabilities through self-avoiding paths with no repeated nodes.
@@
In the graph-based model of stergiopoulos2022automatic, the edge weights indicate potential risks. As a result, an edge with the highest weight represents the worst-case scenario. Depending on the specific problem, the optimal edge or graph can be identified based on the maximum or minimum weight. stergiopoulos2022automatic Another research ammann2005host introduced a host-host model to assess network vulnerabilities and the attack chaining process. This model adopts the point of view of a penetration test to identify the maximum penetration level possible on a host.
@@
In noel2009proactive authors consider proactive intrusion prevention and response using attack graphs, dividing them into unconstrained or constrained graphs. Imposing constraints leads to a reduced scope, thus simplifying complexity and improving understanding of potential threats. An unconstrained graph includes all conceivable attack paths without predefined origins or destinations, presuming that the threat source is unidentified and no targeted assets are specified. In contrast, in a constrained graph, particular starting points are defined, including only those paths that are accessible from the origins. The graph can also end with specified attack goals, concentrating on the protection of critical network assets or the attack surfaces of services or applications employed by users. Several attack surface metrics ou2011quantitative; stuckman2012comparing; ayrour2018modelling have been developed to assess an attacker’s capability to compromise a system or its assets.
@@
The enterprise networks are multilayer systems that represent interconnected entities of various types, i.e. networks of nodes, edges, and layers of different types of interactions interdonato2020multilayer; hammoud2020multilayer; hong2016towards. To facilitate analysis, multiple methods have been proposed to simplify and contract these networks. The study in interdonato2020multilayer presents a comprehensive taxonomy of existing simplification techniques along with the classification of multilayer network methods.
@@ 3 Method
Using the combined graph, we calculate the probabilities of each attack vector $EN$ present in the attack graph at the network level. That is, for each attack step $Si{\rightarrow}Sj$, the corresponding network-level probability ($p_{N}$) of reaching the target attack state from the source attack state via any non-intersecting path through the network is calculated using the Simple Contagion Algorithm presented in kuikka2022efficiency. The combined probability of a successful exploitation of an attack vector between two states is then calculated as the product of the probability that the attack step passes the underlying network and the probability of the vulnerability being exploited:
@@
In order to measure the importance of different vulnerabilities and services, we use metrics based on the probability, expected impact, and expected cumulative impact of an attack. Our interpretation and calculations of the probabilities of different states in the attack graph are based on the principles presented in wang2008attack. First of all, our model assumes that the probability of a vulnerability being exploited (i.e., exploitability) represents the probability that an attack vector will be exploited during an attack. Secondly, we assume that a vulnerability can only be exploited if the attack has already propagated to the preceding state. Thirdly, our model considers the different paths leading to a state disjunctive, which means that the probability of an attack reaching a state is equal to the probability that at least one of the attack paths leading to the state is traversed during an attack.
@@
$ExpectedImpact=Probability*Impact,$
which is motivated by the calculation of risk as a function of likelihood and severity in stergiopoulos2022automatic and assessment of risk as a combination of likelihood and impact in nist2012article. The impact is a property of the vulnerabilities (attack vectors) between the attack states and not of the attack states themselves. Furthermore, the impact values of the different attack vectors leading to a state may differ from each other. To account for this, we consider the expected impact $E[I_{i}]$ of reaching a state as a sum of the expected impacts of the attack vectors $E[I_{j,i}]$ leading to the state $i$:
@@ 4 Results
In order to make our experimental cyberattack case realistic, we chose a network representing what used to be a real-world infrastructure network, shown in the bottom layer in Figure 2. To simulate possible attacks on the services of this network, we used a realistic attack graph that represents vulnerabilities found in a multi-cloud enterprise network originally presented in stergiopoulos2022automatic. To combine the attack graph with the network graph, as a first step, we group the attack states by the affected services as presented in stergiopoulos2022automatic. For each of the services, we selected a representative node in the network graph. The grouping of vulnerabilities in services and the grouping of services to the selected network nodes is listed in Table 1 and shown in the top layer in Figure 2.
The attack graph, as well as the grouped network nodes, are depicted in Figure 3. It consists of 18 states (start, end and 16 attack states), 9 possible attack paths from start to end, and a total of 14 distinct exploitable vulnerabilities. Each vulnerability is represented by one or several directed edges in the attack graph and has both an impact value and an exploitability value calculated using CVSS. The computational methodology for these values is presented in kuikka2024networkmodelling and follows the guidelines shown in Scoring. The list of vulnerabilities, their impact values, exploitability, and their corresponding attack vectors in the attack graph can be seen in Table 2.
@@
Table 2: Vulnerabilities present in the attack graph. The exploitability and impact values are calculated based on the CVSS scores. Scoring Originally presented in kuikka2024networkmodelling.
@@ 4.2 Mitigation of individual vulnerabilities
Similarly to the $C_{IN}$ case, fixing CVE-2004-0840 mitigates all $C_{OUT}(8)$ and fixing both CVE-2007-4752 and CVE-2009-1535 would eliminate all $C_{OUT}(2)$. Although the former can be directly interpreted on the basis of Figure 7, the latter can not, as the relative changes of $C_{OUT}(2)$ after fixing those two vulnerabilities do not sum up to $-1$. This is due to the fact that while fixing CVE-2007-4752 blocks all attack paths targeting node 2 through node 1 and fixing CVE-2009-1535 blocks those that pass through node 4, preventing the attack from progressing through node 2 requires fixing both of them.
@@ 4.3 Securing individual nodes
Figure 11: Relative change in the cumulative impact per attack state,$CE[I_{S18}]$, after adding monitoring to a network service. The horizontal axis represent the number of the node with added monitoring. Note the reversed vertical axis.
@@ Acknowledgements.
[…]
Figure 13: Visual representation of the utilized metrics and sets used in their equations. Blue states are grouped to node $n$ ($G_{n}$), green and black are the states prior to and after $G_{n}$ on attack paths, respectively. (13(a)) An attack graph. (13(b)) Demonstrated metrics. Coloured lines (node $n$ metrics) and solid colour (state $j$ metrics) indicate the attack vectors, whose impact is included in the corresponding metric. (13(c)) Sets used in the equations of the metrics. Coloured lines indicate the set membership of the attack states. The set’s colour matches the metric it belongs to.

### v2

@@ 1 Introduction
Understanding the nature of cyber threats and protecting systems from attacks has become increasingly critical. Cyberattacks occur when malicious actors compromise the confidentiality, integrity, or availability of computer networks and/or services, by exploiting vulnerabilities therein. undef ; undefa Cyber-attackers typically gain access to computer networks by using a series of exploits, in which each exploit facilitates the next. Once they have an initial entry point, they proceed with lateral movement to spread to the rest of the network and reach other hosts and applications within an organization. undefb This series of exploits constitutes an attack path, while the collection of potential attack paths is an attack graph. undefc ; undefd ; undefe
@@
To improve the planning for network security, one needs first to assess the weaknesses in the system. The research presented in undeff ; undefg introduced a model-based framework and a tool to analyse network security across different architectural scenarios. This framework uses a probabilistic model to represent attack paths and countermeasures, and it enables calculation of the probability of success of the attack based on the specific architectural scenario. A study in undefh investigated strategies for protecting critical resources within enterprise networks, focusing mainly on the use of multiple layers of firewalls. This method deploys a perimeter firewall that separates the corporate network from the Internet, along with internal firewalls for critical subnets.
@@
In the present study, we investigate cyber-attacks using a novel approach that combines network modelling and graphs. undefi ; undefj ; undefk This results in an attack graph that is a visual representation of potential paths an attacker could take to compromise a system. undefl ; undefm ; undefn It describes the structure of attack paths as a directed acyclic graph (DAG) of nodes and links between them. undefo Each node in the attack graph can represent a host, network device, or vulnerability of a network system. undefe ; undefp
@@
There are two commonly used types of attack graphs. The first one is a directed graph in which nodes denote system states related to security, such as file modifications, levels of user privileges, and trust relationships among hosts. The edges represent possible exploits or actions by the perpetrators. The second type involves nodes that denote the preconditions or postconditions of an exploit, with edges showing the effects of those conditions. Although our primary focus is on directed acyclic attack graphs, our model can also be used to include more general cyclic and recurrent graphs, or causal graphs. undefq
@@
An attack surface of a system refers to the various ways an intruder can exploit its weaknesses to gain access, undefl including resources that a perpetrator can access. The more resources an attacker can access, the larger the attack surface, making the system less secure. To evaluate the security of a system based on its attack surface, several metrics have been developed. undefr
@@
In the present study, we introduce a quantitative attack modelling approach to combine vulnerability metrics of an enterprise network by integrating the attack graph and the structure of the communication network. Our probabilistic approach uses exploitability metrics to evaluate the likelihood that individual vulnerabilities will be exploited along the attack paths. undeff ; undefs The model estimates the overall probability that a perpetrator successfully gains specific privileges or executes attacks. These probabilities are determined by merging alternative joining attack paths using probability theory for non-exclusive events. undefq
@@
This approach is novel, as it employs a detailed-level network modelling technique undefq that integrates the communication network and attack graph structures. It enables the construction of a quantitative model that describes cyber threats and attacks in the communication network. The outcomes of the model can be presented at varying levels of detail, incorporating attack graph nodes that represent microservices, communication network nodes like servers and network equipment, as well as user services and applications.
@@
Our approach is based on a framework of multilayer network models undeft ; undefu and allows us to analyse cyber threats and design protective solutions against cyber attacks in enterprise networks. Here, we focus on investigating the infrastructure and communication layer, the cyber attack layer, and the application layer, each of which can be divided into sublayers. With the present approach, we can also evaluate the operation and robustness of the system from the point of view of physical network, protocol layers, cyber incidents, user services, and applications. All of these layers have the potential to simultaneously exert detrimental impacts on system users. Moreover, we can construct an extra layer for each user service to evaluate metrics related to both incoming and outgoing effects. This in turn allows us to analyse cyber threats targeting specific services or applications as well as the way an attack propagates within a service between microservices. Different services or application servers provide different options for detection, which affects the likelihood of successful mitigation. It is an essential feature of our method that individual attack paths and attack graphs can be extended to different services and servers in the network topology.
@@ 2 Related Work
In this section, we present a brief overview of related research on modelling cyberattacks and analysing their impact, risk, and resilience in enterprise networks. Here, we focus specifically on graph-based models, attack propagation models, and multilayer network models as well as the metrics to analyse network security. undefv ; undefw ; undefx In undefw the authors present a probabilistic metric for network security by aggregating individual vulnerability measures derived from the scores of the Common Vulnerability Scoring System undefy ; undefs (CVSS) in total network score. The survey on system security metrics presented in undefx proposes a framework to measure system-level security by using the following four sub-metrics: system vulnerabilities, defense capabilities, attack severity, and situational factors. Here the authors introduce a hierarchical ontology to evaluate relationships among these sub-metrics. They also review existing security metrics and highlight their advantages and limitations to guide the development of more effective security metrics. The study in undefz introduces a set of normalised metrics for comparisons across enterprises. It categorises related metrics into four families (victimization, network size, containment, and topology) that aggregate individual scores into family scores. These metrics are presented through interactive visualisations that show trends over time at individual, family, and overall network levels.
@@
In order to analyse a particular part of the network resilience, we need to apply network modelling techniques and graph models to characterise cyber attacks. undefi ; undefj The probability formulas that are used to describe connectivity are also useful in explaining the concept of resilience. undefaa ; undefab Here, the connectivity between two nodes is defined as the probability that there is a functioning connection between them. If the connectivity values between all pairs of neighbouring nodes in the network are known, the connectivity probabilities between any pair of nodes in the network can be calculated. undefaa This calculation can be performed using either simulation methods or algorithmic techniques. An alternative method to this was introduced in undefac , where probabilities were calculated based on spreading probabilities through self-avoiding paths with no repeated nodes.
@@
In the graph-based model of undefm , the edge weights indicate potential risks. As a result, an edge with the highest weight represents the worst-case scenario. Depending on the specific problem, the optimal edge or graph can be identified based on the maximum or minimum weight. undefm Another research undefad introduced a host-host model to assess network vulnerabilities and the attack chaining process. This model adopts the point of view of a penetration test to identify the maximum penetration level possible on a host.
@@
In undefae authors consider proactive intrusion prevention and response using attack graphs, dividing them into unconstrained or constrained graphs. Imposing constraints leads to a reduced scope, thus simplifying complexity and improving understanding of potential threats. An unconstrained graph includes all conceivable attack paths without predefined origins or destinations, presuming that the threat source is unidentified and no targeted assets are specified. In contrast, in a constrained graph, particular starting points are defined, including only those paths that are accessible from the origins. The graph can also end with specified attack goals, concentrating on the protection of critical network assets or the attack surfaces of services or applications employed by users. Several attack surface metrics undefl ; undefaf ; undefag have been developed to assess an attacker’s capability to compromise a system or its assets.
@@
The enterprise networks are multilayer systems that represent interconnected entities of various types, i.e. networks of nodes, edges, and layers of different types of interactions undefu ; undefah ; undefai . To facilitate analysis, multiple methods have been proposed to simplify and contract these networks. The study in undefu presents a comprehensive taxonomy of existing simplification techniques along with the classification of multilayer network methods.
@@ 3 Method
Using the combined graph, we calculate the probabilities of each attack vector $EN$ present in the attack graph at the network level. That is, for each attack step $Si{\rightarrow}Sj$, the corresponding network-level probability ($p_{N}$) of reaching the target attack state from the source attack state via any non-intersecting path through the network is calculated using the Simple Contagion Algorithm presented in undefaj . The combined probability of a successful exploitation of an attack vector between two states is then calculated as the product of the probability that the attack step passes the underlying network and the probability of the vulnerability being exploited:
@@
In order to measure the importance of different vulnerabilities and services, we use metrics based on the probability, expected impact, and expected cumulative impact of an attack. Our interpretation and calculations of the probabilities of different states in the attack graph are based on the principles presented in undefs . First of all, our model assumes that the probability of a vulnerability being exploited (i.e., exploitability) represents the probability that an attack vector will be exploited during an attack. Secondly, we assume that a vulnerability can only be exploited if the attack has already propagated to the preceding state. Thirdly, our model considers the different paths leading to a state disjunctive, which means that the probability of an attack reaching a state is equal to the probability that at least one of the attack paths leading to the state is traversed during an attack.
@@
ExpectedImpact=ProbabilityImpact,ExpectedImpact=Probability*Impact,
which is motivated by the calculation of risk as a function of likelihood and severity in undefm and assessment of risk as a combination of likelihood and impact in undefak . The impact is a property of the vulnerabilities (attack vectors) between the attack states and not of the attack states themselves. Furthermore, the impact values of the different attack vectors leading to a state may differ from each other. To account for this, we consider the expected impact $E[I_{i}]$ of reaching a state as a sum of the expected impacts of the attack vectors $E[I_{j,i}]$ leading to the state $i$:
@@ 4 Results
In order to make our experimental cyberattack case realistic, we chose a network representing what used to be a real-world infrastructure network, shown in the bottom layer in Figure 2. To simulate possible attacks on the services of this network, we used a realistic attack graph that represents vulnerabilities found in a multi-cloud enterprise network originally presented in undefm . To combine the attack graph with the network graph, as a first step, we group the attack states by the affected services as presented in undefm . For each of the services, we selected a representative node in the network graph. The grouping of vulnerabilities in services and the grouping of services to the selected network nodes is listed in Table 1 and shown in the top layer in Figure 2.
The attack graph, as well as the grouped network nodes, are depicted in Figure 3. It consists of 18 states (start, end and 16 attack states), 9 possible attack paths from start to end, and a total of 14 distinct exploitable vulnerabilities. Each vulnerability is represented by one or several directed edges in the attack graph and has both an impact value and an exploitability value calculated using CVSS. The computational methodology for these values is presented in undefal and follows the guidelines shown in undefy . The list of vulnerabilities, their impact values, exploitability, and their corresponding attack vectors in the attack graph can be seen in Table 2.
@@
Table 2: Vulnerabilities present in the attack graph. The exploitability and impact values are calculated based on the CVSS scores. undefy Originally presented in undefal .
@@ 4.2 Mitigation of individual vulnerabilities
Similarly to the $C_{IN}$ case, fixing CVE-2004-0840 mitigates all $C_{OUT}(8)$ and fixing both CVE-2007-4752 and CVE-2009-1535 would eliminate all $C_{OUT}(2)$. Although the former can be directly interpreted on the basis of Figure 7, the latter can not, as the relative changes of $C_{OUT}(2)$ after fixing those two vulnerabilities do not sum up to 1-1. This is due to the fact that while fixing CVE-2007-4752 blocks all attack paths targeting node 2 through node 1 and fixing CVE-2009-1535 blocks those that pass through node 4, preventing the attack from progressing through node 2 requires fixing both of them.
@@ 4.3 Securing individual nodes
[…]